# DexVLA: Vision-Language Model with Plug-In Diffusion Expert for General Robot Control

### v1

@@ Abstract
Enabling robots to perform diverse tasks across varied environments is a central challenge in robot learning. While vision-language-action (VLA) models have shown promise for generalizable robot skills, realizing their full potential requires addressing limitations in action representation and efficient training. Current VLA models often focus on scaling the vision-language model (VLM) component, while the action space representation remains a critical bottleneck. This paper introduces DexVLA, a novel framework designed to enhance the efficiency and generalization capabilities of VLAs for complex, long-horizon tasks across diverse robot embodiments. DexVLA features a novel diffusion-based action expert, scaled to one billion parameters, designed for cross-embodiment learning. A novel embodiment curriculum learning strategy facilitates efficient training: (1) pre-training the diffusion expert that is separable from the VLA on cross-embodiment data, (2) aligning the VLA model to specific embodiments, and (3) post-training for rapid adaptation to new tasks. We conduct comprehensive experiments across multiple embodiments, including single-arm, bimanual, and dexterous hand, demonstrating DexVLA’s adaptability to challenging tasks without task-specific adaptation, its ability to learn dexterous skills on novel embodiments with limited data, and its capacity to complete complex, long-horizon tasks using only direct language prompting, such as laundry folding. In all settings, our method demonstrates superior performance compared to state-of-the-art models like Octo, OpenVLA, and Diffusion Policy.
I Introduction
@@
Enabling robots to perform diverse tasks across varied environments is a central challenge in robotics. Achieving versatility—the ability to solve a variety of tasks across diverse environments, while adapting to language commands, environmental constraints, and unexpected disruptions—is even more demanding. Imitation learning [26, 40, 24, 59, 58, 35], particularly through vision-language-action (VLA) models (e.g., [5, 7, 23, 20]), has shown promise in enabling generalizable skills.
@@
However, realizing the vision of omnipotent robot foundation models faces persistent challenges. Two key bottlenecks hinder progress: 1) Data scarcity: State-of-the-art models, like OpenVLA [23]/Octo [30], rely on massive datasets like the Open-X Embodiment dataset (4,000 hours) [31] or even larger corpora like the 10,000-hour dataset used by $\pi_{0}$ [5]. Collecting such data through human demonstrations is extremely costly and labor-intensive [13, 18, 51, 53]. (2) Architectural imbalance: current VLA models often prioritize scaling the vision-language model (VLM) component, i.e., OpenVLA uses 7B VLM and $\pi_{0}$ uses 3B VLM. Despite its enhanced visual and linguistic understanding through internet-scale data pretraining, the VLM component remains disconnected from the embodied, sensorimotor context of robotic action.
@@
This paper introduces plug-in Diffusion Expert for Vision-Language-Action models, namely DexVLA, a novel framework designed to enhance the efficiency and generalization capabilities of VLA for complex, long-horizon tasks across diverse robot embodiments. We achieve this through two key innovations:
@@
We evaluate DexVLA across diverse embodiments, including single-arm, bimanual, dexterous hand, and mobile bimanual robots, demonstrating its effectiveness on a variety of tasks. DexVLA achieves high success rates on many tasks without task-specific adaptation. For example, it achieves a near full score (0.92) in folding flattened shirts. It can also learn dexterous skills on novel embodiments with fewer than 100 demonstrations, such as pouring drinks with a dexterous hand and packing on a bimanual UR5e. Furthermore, our approach can complete complex, long-horizon tasks like laundry folding using only direct instructions, eliminating the need for a high-level policy model to decompose the task. Importantly, our model is pre-trained on only 100 hours of demonstration data and runs at 60Hz on a single Nvidia A6000 GPU, enabling cost-efficient training and fast inference.
In summary, our contributions are three-fold:
  • We introduce Vision-Language Diffusion Expert (DexVLA), a novel architecture that scales the action expert to facilitate cross-embodiment learning.
  • We present an embodied curriculum learning strategy tailored for efficient and effective training of DexVLA. Furthermore, we enhance the VLA model’s learning by incorporating demonstrations annotated with sub-step reasoning.
  • We conduct comprehensive experiments across multiple embodiments, demonstrating DexVLA’s adaptability to challenging tasks without task-specific adaptation, its ability to learn dexterous skills on novel embodiments with limited data, and its capacity to complete complex, long-horizon tasks using only direct language prompting.
II Related Work
@@
Diffusion models. Diffusion models [11, 32, 19] have emerged as the dominant approach in visual generation. The Diffusion Policy [8] successfully applies the diffusion model to robot learning, demonstrating its ability to model multimodal action distributions. Subsequent research has further developed the Diffusion Policy [61, 48, 34, 38, 41, 42, 3, 4, 10, 25, 37, 47, 29] by applying it to 3D environments [22, 57, 56, 54], scaling its capabilities [63], improving its efficiency [21, 34], and incorporating architectural innovations. There are a number of works investigating the usage of diffusion VLA [50, 5, 49]. Although existing models achieve strong performance and generalization on diverse tasks, they predominantly rely on the capabilities of pre-trained vision-language models. This work proposes a paradigm shift towards the diffusion module, demonstrating that a newly designed diffusion-based action expert, coupled with a novel training strategy, enables VLA models to learn from data more efficiently and effectively.
@@
III Method
This section gives a detailed introduction to our model architecture and training strategy. The overview is presented in Figure 1.
III-A Model Architecture
@@
Our DexVLA model is primarily based on a transformer language model backbone. Following the common framework of VLM models, we employ image encoders to project the robot’s image observations into the same embedding space as the language tokens. For multiple camera views, these visual tokens are concatenated. The VLM component generates two outputs: reasoning tokens and action tokens. The action tokens are passed through a projection module, consisting of two linear layers with LayerNorm. This module is analogous to the connectors designed in vision-language models like LLaVA [28], and serves to transform the VLM’s embedding space to align with the input requirements of the action expert. The reasoning tokens are injected into the policy model using FiLM layers, which scale and shift the parameters of the projection layers within the policy. Consequently, the model can autonomously generate reasoning and leverage this reasoning within the diffusion expert to guide action generation.
Following the standard late fusion VLM recipe [27, 28], image encoders embed the robot’s image observations into the same embedding space as language tokens. We further augment this backbone with robotics-specific inputs and outputs — namely, proprioceptive state and robot actions.
@@
Building diffusion expert. Since action experts dominate the learning process of robot’s action, it is essential to design good neural architecture for better visuomotor policy learning. We utilized the Scale Diffusion Policy (ScaleDP [63]), a variant of the Diffusion Policy in Transformer architecture, where the largest version of ScaleDP is up to 1B parameters. However, the naive ScaleDP is not designed for cross-embodiement pre-training. We design a multi-head output to enable pre-training on ScaleDP with various robot configurations. Each head is responsible for a single robot configuration. This setup is similar to Octo.
@@
Training objectives. Given a batch of input sequences, the overall training loss is defined as a weighted combination of the diffusion loss ($L_{diff}$) and the next-token prediction (NTP) loss ($L_{ntp}$):
$L=L_{diff}+\alpha L_{ntp},$ (1)
where $\alpha$ is a hyperparameter controlling the relative weight of the NTP loss term, $L_{ntp}$. Empirically, without the stage 1 training, the magnitude of $L_{ntp}$ is consistently an order of magnitude smaller than that of $L_{diff}$. Yet, after we adopted stage 1 training to warm up the policy learning ability of diffusion experts, we found that the magnitude of $L_{ntp}$ becomes slightly smaller than that of $L_{diff}$. Therefore, we set $\alpha=1$ in all experiments. This adjustment ensures that both the diffusion and next-token prediction objectives are comparably weighted during training, thus enabling the model to learn effectively from both action prediction and language understanding tasks. The following section provides a comprehensive introduction to the training strategy, detailing how it leverages the capabilities of our model architecture.
III-B Embodied Curriculum Learning
Curriculum learning is an training strategy where a system learns tasks in a progression from simple to complex, mirroring how humans acquire skills. Our three-stage training strategy implements an embodied curriculum, where the agent first learns generalizable motor skills from cross-embodiment data (Stage 1), then adapts to its specific physical form (Stage 2), and finally refines task-specific behaviors (Stage 3). This mirrors human skill acquisition, where foundational abilities (e.g., grasping) precede specialized expertise (e.g., folding clothes). By decoupling embodiment-agnostic and embodiment-specific learning, we reduce data needs by 60% compared to end-to-end training.
A well-designed training strategy is critical for optimizing deep neural networks. Approaches that align with a network’s inherent training dynamics ensure more efficient and effective data utilization. Our framework, DexVLA, targets general robotic control by integrating a Vision-Language Model (VLM) with a diffusion expert. Leveraging its modular architecture—which combines two distinct components—we propose a three-stage training strategy that systematically addresses: (1) learning dexterous manipulation skills to enable the model to complete complex tasks; and (2) cross-embodiment learning to adapt the model to diverse robotic platforms.
III-B1 Stage 1: Cross-Embodiment Pre-training
The vision-language-action models can be viewed as a composite of two distinct components. At the top of the architecture lies the vision-language model (VLM), which processes both visual input and language instructions, mapping them into a shared embedding space. This shared space is pre-trained using internet-scale data, enabling a wide range of capabilities, including language understanding, multimodal understanding, and various other vision-text tasks. However, despite its extensive training, the VLM lacks the ability to physically interact with diverse objects in real-world environments.
To effectively pre-train the action expert, we leverage all available data while temporarily decoupling it from the VLM component. This allows us to focus on developing a robust action generation capability independent of language grounding. While CNN architectures, such as ResNet, are commonly used as visual encoders, the visual encoders within VLM models typically employ a Transformer architecture. Therefore, for ease of alignment in subsequent stages, we utilize a ViT architecture for visual encoding. For language embedding extraction, we use DistilBERT [39]. The resulting language embeddings are then integrated into the model using FiLM layers, consistent with previous work. A key insight for pre-training the diffusion expert is the necessity of decomposing long-horizon tasks (e.g., bussing tables, laundry folding) into sub-tasks. These tasks, often spanning beyond 2 minutes, prove challenging for the diffusion expert to learn effectively from a single language instruction. Therefore, we annotate sub-step instruction within these long-horizon tasks to provide a more structured learning signal. Pre-training with sub-steps is crucial for strong performance. We demonstrate the importance of sub-step reasoning in the ablation study (see Appendix). Our empirical observations show that VLA trained without this pre-training frequently skip critical steps in very long tasks. Sub-step annotations are typically provided every five seconds of the demonstration.
Pre-training the large-scale robotic data on the diffusion-based action expert, rather than the entire VLA, offers several advantages. First, our diffusion expert has a significantly smaller parameter count compared to VLAs (1 billion vs. 3 billion), resulting in much faster training. Quantitatively, pre-training on the diffusion expert is five times faster than pre-training on the full VLA model.
@@ III-B1 Stage 1: Cross-Embodiment Pre-training
III-B2 Stage 2: Embodied-Specific Alignment
While Stage 1 learns basic motor skills from cross-embodied data, this cross-embodiment learning can potentially compromise performance on the target embodiment, making it unsuitable for real-world deployment. Stage 2 addresses this by training the model with embodiment-specific data, aligning the abstract vision-language representations from the VLM with the diffusion expert. Therefore, we filter the dataset to include only embodiment-specific data, ensuring each sample involves a single embodiment. Mirroring techniques employed in vision-language models like LLaVA [28, 27], this stage focuses on aligning the target embodiment’s action space with its corresponding camera views and accompanying language instructions. Specifically, we jointly train the VLM model, the projection layer, and the diffusion expert on this embodiment-specific data, while freezing the VLM’s visual encoder. This joint training allows the diffusion expert to effectively ground the high-level vision-language understanding from the VLM in the specific motor control space of the target robot. Following Stage 2 training, we observe that the model exhibits proficiency in performing a range of tasks on the target embodiment, such as shirt folding, demonstrating the effectiveness of the embodiment-specific training.
@@ III-B2 Stage 2: Embodied-Specific Alignment
III-B3 Stage 3: Task-Specific Adaptation
This phase refines the model’s ability to execute downstream tasks skillfully and fluently, analogous to the post-training stage in large language models where the model is fine-tuned on domain-specific data. For simpler, less generalization-dependent tasks, such as shirt folding, table bussing, or bin picking with trained objects, task-specific training is unnecessary as the model already performs well. However, complex, dexterity-demanding tasks require the model to learn fine-grained, context-dependent actions. Therefore, effective post-training relies on a high-quality dataset of expert demonstrations exhibiting consistent and fluent task execution strategies focused on behaviors that promote successful task completion. It is worth noting that we utilize sub-step annotated language data in both Stage 2 and Stage 3. However, instead of directly using these sub-step reasoning as instructional input, we employ them as intermediate language output, compelling the model to learn and generate these sub-step language descriptions. This approach has proven highly effective, enabling our model to perform complex, long-horizon tasks such as laundry folding. While other VLAs, like $\pi_{0}$ [5], can also complete such tasks, they rely on high-level policy models like SayCan [1] to identify the task state and provide next-step instructions. In contrast, our framework leverages the VLM backbone as an implicit high-level policy model. This allows the model to internally interpret the task’s state and inject this understanding into the policy to guide action generation, eliminating the need for an external high-level policy module.
III-C Implementation Details
In our experiments, we employ the Qwen-2-VL 2B [46] model as our base vision-language model. We utilize up to three camera views. Our diffusion expert comprises 1 billion parameters. Two MLP layers, each followed by LayerNorm, connect the vision-language model to the action experts. We use FiLM to inject the final embedding from the vision-language model into the diffusion expert, applying it to each of the two MLPs. During Stage 1 training, we use our collected dataset. Stage 2 training is restricted to data sharing the same embodiment. Post-training is performed selectively, only on tasks where it is explicitly mentioned. Our pretrained dataset has a total of 100 hours of data with 91 different tasks. The detailed data allocation is referred to the Figure 4. For clarity, we report data quantity in hours rather than trajectories, as the duration of tasks varies significantly (i.e., laundry folding takes considerably longer than other tasks). This method of data quantification is consistent with that used by $\pi_{0}$. The detailed hyperparameters are listed in Table I. Specific details on implementation can be found in the Appendix.
TABLE I: Detailed hyperparameters for training Dex-VLA.
Hyperparameters Stage 1 Stage 2 Stage 3
Learning rate 1e-4 2e-5 2e-5
LR scheduler Constant Constant Cosine
Weight decay 0.0 0.0 0.0
Optimizer AdamW($\beta_{1}=0.9$,$\beta_{2}$=0.95)
Training epochs 5 5 5
Data Cross-embodied Embodied-specific Task-specific
Figure 4: Overview of our dataset for stage 1 training. Our dataset comprises approximately 100 hours of collected data spanning 91 distinct tasks. The majority of this data was collected using two robot platforms: the Agilex (ARX arm) (42.7%) and the single Franka Emika robot (34.7%). The “ARX arm” and “PIPER arm” represent two distinct robotic arm configurations, both featuring six degrees of freedom (6-DoF) but differing in their kinematic structures and operational characteristics.
III-D Robot System Setup
Our evaluation is conducted on four different robot configurations across 10 tasks. These setups are summarized in the following list and visualized in the accompanying figure:
IV Real World Experiments
@@
Figure 5: Examples of tasks without task-specific adaptation. We assessed our model’s performance after stage 2 training using three tasks: bin-picking easy (top), shirt folding (middle), and table bussing easy (bottom). Successful execution of these tasks necessitates a combination of dexterous manipulation skills.

### v2

@@ Abstract
Enabling robots to perform diverse tasks across varied environments is a central challenge in robot learning. While vision-language-action (VLA) models have shown promise for generalizable robot skills, realizing their full potential requires addressing limitations in action representation and efficient training. Current VLA models often focus on scaling the vision-language model (VLM) component, while the action space representation remains a critical bottleneck. This paper introduces DexVLA, a novel framework designed to enhance the efficiency and generalization capabilities of VLAs for complex, long-horizon tasks across diverse robot embodiments. DexVLA features a novel diffusion-based action expert, scaled to one billion parameters, designed for cross-embodiment learning. A novel embodiment curriculum learning strategy facilitates efficient training: (1) pre-training the diffusion expert on cross-embodiment data, (2) aligning the VLA model to specific embodiments, and (3) post-training for rapid adaptation to new tasks. We conduct comprehensive experiments across multiple embodiments, including single-arm, bimanual, and dexterous hand, demonstrating DexVLA’s adaptability to challenging tasks without task-specific adaptation, its ability to learn dexterous skills on novel embodiments with limited data, and its capacity to complete complex, long-horizon tasks using only direct language prompting, such as laundry folding. In all settings, our method demonstrates superior performance compared to state-of-the-art models like OpenVLA and $\pi_{0}$.
Keywords: Vision-Language-Action Model, Robotic Manipulation
footnotetext: : denotes equal contribution. denotes corresponding author.
1Midea Group, 2East China Normal University, 3Shanghai University
1 Introduction
@@
Enabling robots to perform diverse tasks across varied environments is a central challenge in robotics. Achieving versatility—the ability to solve a variety of tasks across diverse environments, while adapting to language commands, environmental constraints, and unexpected disruptions—is even more demanding. Imitation learning [1, 2, 3, 4, 5, 6, 7], particularly through vision-language-action (VLA) models [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], has shown promise in enabling generalizable skills.
@@
However, realizing the vision of omnipotent robot foundation models faces persistent challenges. Two key bottlenecks hinder progress: 1) Data scarcity: State-of-the-art models, like OpenVLA [20] and Octo [21], rely on massive datasets like the Open-X Embodiment dataset (4,000 hours) [22] or even larger corpora like the 10,000-hour dataset used by $\pi_{0}$ $\pi_{0.5}$ [8, 13]. Collecting such data through human demonstrations is extremely costly and labor-intensive [23, 24, 25, 26]. (2) Architectural imbalance: current VLA models often prioritize scaling the vision-language model (VLM) component, i.e., OpenVLA uses 7B VLM and $\pi_{0}$ uses 3B VLM. Despite its enhanced visual and linguistic understanding through internet-scale data pretraining, the VLM component remains disconnected from the embodied, sensorimotor context of robotic action.
@@
This paper introduces Plug-in Diffusion Expert for Vision-Language-Action models, namely DexVLA, a novel framework designed to enhance the data efficiency and generalization capabilities of VLA for complex, long-horizon tasks across diverse robot embodiments. We achieve this through two key innovations:
@@
We evaluate DexVLA across diverse embodiments, including single-arm, bimanual, dexterous hand, and mobile bimanual robots, demonstrating its effectiveness on a variety of tasks. DexVLA achieves high success rates on many tasks without task-specific adaptation. For example, it achieves a near full score in folding flattened shirts. It can also learn dexterous skills on novel embodiments with fewer than 100 demonstrations, such as pouring drinks with a dexterous hand and packing on a bimanual robot. Furthermore, when directly prompting VLA model on completing complex, long-horizon tasks like laundry folding, DexVLA outperforms $\pi_{0}$ by a large margin. Importantly, our model is pre-trained on only 100 hours of demonstration data and runs at 60Hz on a single Nvidia A6000 GPU, enabling cost-efficient training and fast inference.
2 Related Work
@@
Diffusion models. Diffusion models [48, 49, 50] have emerged as the dominant approach in visual generation. The Diffusion Policy [30] successfully applies the diffusion model to robot learning, demonstrating its ability to model multimodal action distributions. Subsequent research has further developed the Diffusion Policy [51, 52, 53, 7, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63] by applying it to 3D environments [64, 65, 66, 67], scaling its capabilities [29], improving its efficiency [68, 53], and incorporating architectural innovations. There are a number of works investigating the usage of diffusion VLA [43, 8, 69]. Although existing models achieve strong performance and generalization on diverse tasks, they predominantly rely on the capabilities of pre-trained vision-language models. This work proposes a paradigm shift towards the diffusion module, demonstrating that a newly designed diffusion-based action expert, coupled with a novel training strategy, enables VLA models to learn from data more efficiently and effectively.
@@
3 Method
3.1 Model Architecture
@@
Our DexVLA model is primarily based on a transformer language model backbone. We use Qwen2-VL [27] as the base VLM model. Following the common framework of VLM models, we employ image encoders to project the robot’s image observations into the same embedding space as the language tokens. For multiple camera views, these visual tokens are concatenated. The VLM component generates two outputs: reasoning tokens and action tokens. The action tokens are passed through a projection module, consisting of two linear layers with LayerNorm. This module is analogous to the connectors designed in vision-language models like LLaVA [28], and serves to transform the VLM’s embedding space to align with the input requirements of the action expert. The reasoning tokens are injected into the policy model using FiLM layers, which scale and shift the parameters of the projection layers within the policy. Consequently, the model can autonomously generate reasoning and leverage this reasoning within the diffusion expert to guide action generation. The overview is presented in Figure 1.
@@
Building diffusion expert. Since action experts dominate the learning process of robot’s action, it is essential to design a good neural architecture for better visuomotor policy learning. We utilized the Scale Diffusion Policy (ScaleDP [29]), a variant of the Diffusion Policy in Transformer architecture, where the largest version of ScaleDP is up to 1B parameters. However, the naive ScaleDP is not designed for cross-embodiment pre-training. We put a multi-head output to enable pre-training on ScaleDP with various robot configurations. Each head is responsible for a single robot configuration. This setup is similar to Octo [21].
@@
Training objectives. Given a batch of input sequences, the overall training loss is defined as a weighted combination of the diffusion loss ($L_{diff}$) and the next-token prediction loss ($L_{ntp}$), which are $L=L_{diff}+\alpha L_{ntp}$. For all experiments, we set $\alpha=1$, as we observe that $L_{ntp}$ converges during the early stages of training. Consequently, this setup allows the model to primarily focus on learning robot action prediction based on reasoning and instructions.
@@
Figure 4: Detailed hyperparameters for training DexVLA. Hyperparameters Stage 1 Stage 2 Stage 3 Learning rate 1e-4 2e-5 2e-5 LR scheduler Constant Constant Cosine Weight decay 0.0 0.0 0.0 Optimizer AdamW($\beta_{1}=0.9$,$\beta_{2}$=0.95) Training epochs 5 5 5 Data Cross-embodied Embodied-specific Task-specific
3.2 Embodied Curriculum Learning
Curriculum learning is an training strategy where a system learns tasks in a progression from simple to complex, mirroring how humans acquire skills. Our three-stage training strategy implements an embodied curriculum, where the policy network first learns generalizable motor skills from cross-embodiment data (Stage 1), then adapts to its specific physical form (Stage 2), and finally refines task-specific behaviors (Stage 3). This mirrors human skill acquisition, where foundational abilities (e.g., grasping) precede specialized expertise (e.g., folding clothes).
A well-designed training strategy is critical for optimizing deep neural networks. Approaches that align with a network’s inherent training dynamics ensure more efficient and effective data utilization. The  DexVLAtargets general robotic control by integrating a VLM with a diffusion expert. Leveraging its modular architecture—which combines two distinct components—we propose a three-stage training strategy that systematically addresses: (1) learning dexterous manipulation skills to enable the model to complete complex tasks; and (2) cross-embodiment learning to adapt the model to diverse robotic platforms.
Stage 1: Cross-embodiment pre-training. The vision-language-action models can be viewed as a composite of two distinct components. At the top of the architecture lies the vision-language model (VLM), which processes both visual input and language instructions, mapping them into a shared embedding space. This shared space is pre-trained using internet-scale data, enabling a wide range of capabilities, including language understanding, multimodal understanding, and various other vision-text tasks. However, despite its extensive training, the VLM lacks the ability to physically interact with diverse objects in real-world environments.
To effectively pre-train the action expert, we leverage all available data while temporarily decoupling it from the VLM component. This allows us to focus on developing a robust action generation capability independent of language grounding. We use a ResNet-50 as image encoders, aligning with DP [30] and DistilBERT [31] as a language embedding model. The resulting language embeddings are then integrated into the model using FiLM layers, consistent with previous work [32, 33].
Stage 2: Embodied-specific alignment. While Stage 1 learns basic motor skills from cross-embodied data, this cross-embodiment learning can potentially compromise performance on the target embodiment, making it unsuitable for real-world deployment. Stage 2 addresses this by training the model with embodiment-specific data, aligning the abstract vision-language representations from the VLM with the diffusion expert. Therefore, we filter the dataset to include only embodiment-specific data, ensuring each sample involves a single embodiment. Mirroring techniques employed in vision-language models like LLaVA [28, 34], this stage focuses on aligning the target embodiment’s action space with its corresponding camera views and accompanying language instructions. Specifically, we jointly train the VLM model, the projection layer, and the diffusion expert on this embodiment-specific data, while freezing the VLM’s visual encoder. This joint training allows the diffusion expert to effectively ground the high-level vision-language understanding from the VLM in the specific motor control space of the target robot. Following Stage 2 training, we observe that the model exhibits proficiency in performing a range of tasks on the target embodiment, such as shirt folding, demonstrating the effectiveness of the embodiment-specific training.
@@ 3.2 Embodied Curriculum Learning
Figure 5: Examples of tasks without task-specific adaptation. We assessed our model’s performance after stage 2 training using three tasks: bin-picking easy (top), shirt folding (middle), and table bussing easy (bottom).